# Elastic displacements and viscous hydrodynamic flows in wedge-shaped geometries featuring a straight edge: Green’s functions for forces oriented parallel to the edge

### v1

@@ V Solution in the limit of a planar boundary
Here, $s$ denotes the distance from the singularity, as defined earlier in Eq. (19), and
@@
representing the distance from the image of the point force relative to the planar interface.
@@ VI Conclusions
The data and numerical codes that support the findings of this study are available from the corresponding author upon reasonable request.
@@ A.5 A few remarks on parity with respect to $z$
For instance, in Ref. 39, the FKL transform of $(z/r)f$ is expressed in terms of the imaginary part when presenting the coefficients $C^{A}$ and $C^{S}$ in their Eq. (3.9), where $f$ is an even function of $z$. In contrast to that, the general formulation in Eq. (79) yields significantly simpler expressions, which have proven highly useful in our calculations.
@@ C.1 Integral $\mathscr{L}_{1}$
$\mathscr{L}_{1}=\frac{1}{\pi}\int_{0}^{\pi}\mathrm{d}t\int_{0}^{\infty}\frac{\operatorname{ch}(a{p})}{\operatorname{ch}(\pi{p})}\,\frac{\mathrm{d}{p}}{\left(\zeta(t)\right)^{{p}+\frac{1}{2}}}\,,$ (95)
@@
Integration with respect to $p$ is performed using the residue method [48] by integrating along a rectangular contour in the lower half-plane. Specifically, we choose an anticlockwise-oriented contour in the complex plane, composed of four rectilinear segments. The vertices are positioned at $(R,0)$, $(R,R)$, $(-R,R)$, and $(-R,0)$. In the limit $R\to\infty$, only the integral along the real axis does not vanish, and it evaluates to $2i\pi$ times the sum of the residues at $z=(2n+1)i/2$ for $n\geq 0$. We obtain
@@
Then, the limits of integration become $[\pi,0]$. Accordingly,
@@ C.2 Integral $\mathscr{L}_{2}$
This integral involves the associated Legendre function of order $-1$ and is related to the evaluation of the polar and azimuthal components of the displacement field. We define
@@
For ${q}=i{p}-\frac{1}{2}$ and $m=-1$, this expression reads
@@
It is worth noting that $\mathscr{L}_{1}$ and $\mathscr{L}_{2}$ diverge for $a$ outside the range $[-\pi,\pi]$. However, since $a=\theta+\beta$ remains within this interval, convergence is ensured.

### v2

@@ V Solution in the limit of a planar boundary
Here,
@@
represents the distance from the image of the point force relative to the planar interface.
@@ VI Conclusions
Beyond the examples already mentioned, the derived Green’s function for a wedge-shaped geometry opens up a range of further applications. In microfluidic systems, wedge-like features commonly appear in channel intersections or patterned substrates; here, the Green’s function can aid in modeling electrokinetic flows under applied electric fields. Similarly, in acoustofluidic contexts, it may help describe acoustic streaming effects and wave propagation in confined viscous media. The hydrodynamic interactions of colloidal particles or microswimmers near wedge boundaries can also be quantified, shedding light on their aggregation or diffusive behaviors. The Green’s function further supports the design of edge-guided actuation in soft robotics, where embedded stimuli in a wedge-like structure can direct movement or amplify local responses. Finally, in biological systems, many tissues feature sharp anatomical boundaries—such as near joints or organ interfaces—and the Green’s function may contribute to modeling localized deformations due to internal or external forces in these complex geometries.
We remark that the geometry of the wedge-shaped geometry itself has no intrinsic length scale. This makes our solution rather general in application, as long as our assumed conditions such as linear elasticity and low-Reynolds-number flows are met. The relative position of applying the force within the wedge enters the expression of the Green’s function.
@@
Mathematica scripts used to generate the data files are included in the Supporting Information.
@@ A.5 A few remarks on parity with respect to $z$
For instance, in Ref. 27, the FKL transform of $(z/r)f$ is expressed in terms of the imaginary part when presenting the coefficients $C^{A}$ and $C^{S}$ in their Eq. (3.9), where $f$ is an even function of $z$. In contrast to that, the general formulation in Eq. (79) yields significantly simpler expressions, which have proven highly useful in our calculations.
@@ C.1 Integral $\mathscr{L}_{1}$
$\mathscr{L}_{1}=\frac{1}{\pi}\int_{0}^{\pi}\mathrm{d}t\int_{0}^{\infty}\frac{\operatorname{ch}(a{p})}{\operatorname{ch}(\pi{p})}\,\frac{\mathrm{d}{p}}{\left(\zeta(t)\right)^{i{p}+\frac{1}{2}}}\,$, (95)
@@
Integration with respect to $p$ is performed using the residue method [73] by integrating along a rectangular contour in the lower half-plane. Specifically, we choose an clockwise-oriented contour in the complex plane, composed of four rectilinear segments. The vertices are positioned at (R,0)(-R,0), $(R,0)$, (R,R)(R,-R), and (R,R)(-R,-R). In the limit $R\to\infty$, only the integral along the real axis does not vanish, and it evaluates to $2i\pi$ times the sum of the residues at z=(2n+1)i/2z=-(2n+1)i/2 for $n\geq 0$. We obtain
@@
Then, the limits of integration become $\pi,0]$. Accordingly,
@@ C.2 Integral $\mathscr{L}_{2}$
This integral involves the associated Legendre function of order 1-1 and is related to the evaluation of the polar and azimuthal components of the displacement field. We define
@@
For ${q}=i{p}-\frac{1}{2}$ and m=1m=-1, this expression reads
@@
It is worth noting that $\mathscr{L}_{1}$ and $\mathscr{L}_{2}$ diverge for $a$ outside the range $\pi,\pi]$. However, since $a=\theta+\beta$ remains within this interval, convergence is ensured.
Supporting Information
The Mathematica files FS-FS-RThe-Plane, NS-NS-RThe-Plane, and NS-FS-RThe-Plane generate output data for the displacement field in the radial-azimuthal plane, corresponding to the 2FRS, 2NOS, and NOS–FRS boundary conditions, respectively. The files FS-FS-RZ-Plane, NS-NS-RZ-Plane, and NS-FS-RZ-Plane provide the corresponding results in the radial-axial plane. The MATLAB scripts RThe_Plane_Plotter and RZ_Plane_Plotter read the output from the Mathematica files and plot the resulting streamlines in the radial-azimuthal and radial-axial planes, respectively.